# Satisficing Regret Minimization in Bandits

### v2

@@ C.2 Proof of Theorem 4
Appendix D Choices of $C_{1},\alpha,\beta$ for Several Bandit Problem Classes
@@
In this section, we provide a detailed discussion on the choice of $C_{1},\alpha,\beta$ for several bandit problem classes. In three cases
D.1 Finite-armed Bandits
@@
D.2 Concave Bandits
Consider any concave bandits with $\mathcal{X}\subseteq\mathbb{R}^{d}$ being a compact and convex set contained in a Euclidean ball with radius $1$, and reward function being concave and $1$-Lipschitz. By Theorem 2 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 2 in Agarwal et al. (2011) is upper bounded by
@@ D.2 Concave Bandits
Furthermore, if the instance of concave bandit is one-dimensional, i.e., $\mathcal{X}=[0,1]$, by Theorem 1 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 1 in Agarwal et al. (2011) is bounded by
@@ D.3 Lipschitz Bandits
Using the results in Section D.1, we have $c_{\text{ALG}}=11$, thus the regret of the Uniform UCB algorithm is bounded by
@@
Appendix E Ablation Study for SELECT
In this section, we use an ablation study to demonstrate that all three steps of SELECT plays a vital role in obtaining a constant regret bound. Specifically, we compare the performance of SELECT with the following adapted versions of SELECT:
  • SELECT without step 1: instead of identifying candidate satisficing arms using a learning oracle, the candidate satisficing arm is drawn uniformly at random from the arm set before we proceed to forced sampling and LCB tests in each round;
@@ Appendix E Ablation Study for SELECT
Appendix F Robustness on the choice of $\gamma_{i}$
@@
In Algorithm 1, we define $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$ for all $i\in\mathbb{N}_{+}$, but all our results still hold if all $\gamma_{i}$ are multiplied by a constant. In this section, we conduct numerical experiments on testing the robustness of SELECT in the choice $\gamma_{i}$.

### v3

@@ H.5 Proof of Theorem 10
Appendix I Choices of $C_{1},\alpha,\beta$ for Several Bandit Optimization Problem Classes
@@
In this section, we provide a detailed discussion on the choice of $C_{1},\alpha,\beta$ for several bandit optimization problem classes. In three cases
I.1 Finite-armed Bandits
@@
I.2 Concave Bandits
Consider any concave bandits with $\mathcal{A}\subseteq\mathbb{R}^{d}$ being a compact and convex set contained in a Euclidean ball with radius $1$, and reward function being concave and $1$-Lipschitz. By Theorem 2 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 2 in Agarwal et al. (2011) is upper bounded by
@@ I.2 Concave Bandits
Furthermore, if the instance of concave bandit is one-dimensional, i.e., $\mathcal{A}=[0,1]$, by Theorem 1 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 1 in Agarwal et al. (2011) is bounded by
@@ I.3 Lipschitz Bandits
Using the results in Section I.1, we have $c_{\text{ALG}}=11$, thus the regret of the Uniform UCB algorithm is bounded by
@@
Appendix J Ablation Study for SELECT
In this section, we use an ablation study to demonstrate that all three steps of SELECT play a vital role in obtaining a constant satisficing regret. Specifically, we compare the performance of SELECT with the following adapted versions of SELECT:
  • SELECT without Step 1: instead of identifying candidate satisficing arms using a learning oracle, the candidate satisficing arm is drawn uniformly at random from the arm set before we proceed to forced sampling and LCB test in each round;
@@
Appendix K Robustness on the choice of $\gamma_{i}$
@@
In Algorithm 1, we define $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$ for all $i\in\mathbb{Z}_{+}$, but all our results still hold if all $\gamma_{i}$ are multiplied by a constant. In this section, we conduct numerical experiments on testing the robustness of SELECT in the choice $\gamma_{i}$.